OpenStudy (shamil98):

h(x) = 3x / 1 + √x quotient rule? 3(1+1√x) - (1/2(x^1/2))(3x) / (1+√x)^2 @iambatman

OpenStudy (anonymous):



OpenStudy (shamil98):

yeah

OpenStudy (anonymous):

Yup

OpenStudy (shamil98):

alrighty.

OpenStudy (anonymous):

Set it up :p

OpenStudy (shamil98):

it says to not simplify so that's the final answer.

OpenStudy (shamil98):

3(1+1√x) - (1/2(x^1/2))(3x) / (1+√x)^2

OpenStudy (anonymous):

Did you understand the process?

OpenStudy (shamil98):

Yeah. Just plug in the values. f(x) = 3x f'(x) = 3 g(x) = 1 + √x g'(x) = 1/2√x